# A Guide to Misinformation Detection Datasets

### v1

@@ Abstract
Misinformation is a complex societal issue, and mitigating solutions are difficult to create due to data deficiencies. To address this problem, we have curated the largest collection of (mis)information datasets in the literature, totaling 75. From these, we evaluated the quality of all of the 36 datasets that consist of statements or claims. We assess these datasets to identify those with solid foundations for empirical work and those with flaws that could result in misleading and non-generalizable results, such as insufficient label quality, spurious correlations, or political bias. We further provide state-of-the-art baselines on all these datasets, but show that regardless of label quality, categorical labels may no longer give an accurate evaluation of detection model performance. We discuss alternatives to mitigate this problem. Overall, this guide aims to provide a roadmap for obtaining higher quality data and conducting more effective evaluations, ultimately improving research in misinformation detection. All datasets and other artifacts are available at misinfo-datasets.complexdatalab.com.
1 Introduction
@@
Misinformation is a pressing concern for society, already causing significant negative impacts and posing even greater risks with the advent of generative AI (Torkington, 2024). Extensive research has been devoted to this problem, yet it remains unresolved. There has been considerable recent progress in methods, especially leveraging LLMs (Chen & Shu, 2023). However, to fuel further progress, we also need strong and reliable data.
@@
Multiple studies have identified data availability, and especially data quality, as a barrier in this domain. To begin with, obtaining high quality veracity labels is challenging and time-consuming, even for experts (Zubiaga et al., 2016). Shortcuts, though, can cause severe spurious correlations (Pelrine et al., 2021; Wu & Hooi, 2022), and even with high quality labels there can be issues with ambiguity of input texts (Pelrine et al., 2023). There are many surveys of the method landscape (Shu et al., 2017; Oshikawa et al., 2018; Zhou & Zafarani, 2020; Chen & Shu, 2023), but analysis of datasets has often been limited in scale (Pelrine et al., 2021; Wu & Hooi, 2022; Pelrine et al., 2023) or depth.
@@
In this study, we present the largest scale survey of datasets in the literature to date, curating 75 datasets along with corresponding descriptive analyses and categorizations. This is nearly 3 times as many as other dataset-focused surveys like Hamed et al. (2023), and many times more than general surveys like Ali et al. (2022); Shu et al. (2017); Oshikawa et al. (2018); Zhou & Zafarani (2020). We provide a summary of each dataset, along with key information like topic, size, modality, languages, geographic region, and time period.
Then, we focus on 36 datasets that include claims, which serve as key sources of atomic data for misinformation detection, and we evaluate their quality in depth. First, we assess label quality, examining the labeling process and comparing the advantages and limitations of expert fact-checking, crowdsourced labels, source-based evaluations and algorithmic labeling. Then, we examine three types of potential spurious correlations and bias, that could lead to predictions based on invalid, non-generalizable signals. In particular, we start by looking at keyword based correlations—which can also serve as a proxy for topic or event. Then we look at temporal correlations, and finally the political leaning of the statements. These quality evaluations, combined with our descriptive analysis, offer practical insights for selecting datasets in future misinformation research.
@@
Once datasets are chosen, implementation and evaluation questions remain. To help address these, we first present a unified formatting and labeling schema for all 36 claims datasets. Next, we establish state-of-the-art baselines using GPT-4 with and without web search to collect evidence. Following this analysis, we find that standard evaluation metrics like accuracy and F1, when computed simply in relation to ground truth labels, are no longer sufficient to evaluate leading generative methods for misinformation detection and could lead to invalid conclusions. We provide initial work on an alternative evaluation as a starting point for future research, and suggest such evaluation research is urgently needed in the misinformation detection domain.
In summary, we present a guide to misinformation datasets, including:
  • The largest scale survey of such datasets, with over 75 identified and analyzed, spanning diverse topics and domains such as politics, health, climate, and more.
  • An in depth evaluation of the quality of 36 datasets focused on claims, identifying limitations to take into account when selecting and using them.
  • Practical tools for research using those datasets: a unified file and label schema, and state-of-the-art baselines.
  • Analysis of evaluation suggesting simple metrics like accuracy and F1 may be obsolete in this domain, and initial work on alternatives.

### v2

# A Guide to Misinformation Detection Data and Evaluation

@@ Abstract.
Misinformation is a complex societal issue, and mitigating solutions are difficult to create due to data deficiencies. To address this, we have curated the largest collection of (mis)information datasets in the literature, totaling 75. From these, we evaluated the quality of 36 datasets that consist of statements or claims, as well as the 9 datasets that consist of data in purely paragraph form. We assess these datasets to identify those with solid foundations for empirical work and those with flaws that could result in misleading and non-generalizable results, such as spurious correlations, or examples that are ambiguous or otherwise impossible to assess for veracity. We find the latter issue is particularly severe and affects most datasets in the literature. We further provide state-of-the-art baselines on all these datasets, but show that regardless of label quality, categorical labels may no longer give an accurate evaluation of detection model performance. Finally, we propose and highlight Evaluation Quality Assessment (EQA) as a tool to guide the field toward systemic solutions rather than inadvertently propagating issues in evaluation. Overall, this guide aims to provide a roadmap for higher quality data and better grounded evaluations, ultimately improving research in misinformation detection. All datasets and other artifacts are available at misinfo-datasets.complexdatalab.com.
Misinformation, Datasets, Survey, Quality, Detection
journalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; August 03–07, 2025; Toronto, ON
1. Introduction
@@
Misinformation is a pressing concern for society, already causing significant negative impacts and posing even greater risks with the advent of generative AI (Torkington, 2024). Extensive research has been devoted to this problem, yet it remains unresolved. There has been considerable recent progress in methods, especially leveraging LLMs to detect false information at scale (Chen and Shu, 2023). However, to fuel further progress, we also need strong and reliable data.
@@
Multiple studies have identified data availability, and especially data quality, as a barrier for reliable misinformation detection. To begin, obtaining high quality veracity labels is challenging and time-consuming, even for experts (Zubiaga et al., 2016). Shortcuts, though, can cause severe spurious correlations (Pelrine et al., 2021; Wu and Hooi, 2022), and even with high quality labels there can be issues with ambiguity of input texts (Pelrine et al., 2023). While several surveys have mapped the methodological landscape in this domain (Shu et al., 2017; Oshikawa et al., 2018; Zhou and Zafarani, 2020; Chen and Shu, 2023), the analysis of dataset quality remains either limited in scale (Pelrine et al., 2021; Wu and Hooi, 2022; Pelrine et al., 2023) or lacking in depth.
@@
To overcome this problem, we present the largest-scale survey in the literature to date, curating 75 datasets with comprehensive descriptive analyses and categorizations. This is nearly three times as many as other dataset-focused surveys like (Hamed et al., 2023), and many times more than general surveys like those of (Ali et al., 2022; Shu et al., 2017; Oshikawa et al., 2018; Zhou and Zafarani, 2020). We provide a summary of each dataset, along with key information like topic, size, modality, languages, geographic region, and time period.
We further focus on 36 datasets that include claims and 9 datasets that include paragraphs, and evaluate their quality in depth. First, we examine two types of potential spurious correlations that could lead to predictions based on invalid, non-generalizable signals. In particular, we start by looking at keyword based correlations before following up with temporal correlations. Both of these can serve as proxies for topics, events, and other correlated information. We then assess whether the examples in these datasets are actually feasible to assess for veracity at all. We find that most datasets contain substantial ambiguities and other issues, such that over half of the claims data may be infeasible for methods without evidence retrieval, with a large portion remaining infeasible even with retrieval. These quality assessments offer both immediately practical insights for selecting the most reliable datasets, and significant implications for directions in the field of misinformation research overall.
@@
Our work also addresses challenges of model evaluation after researchers complete the dataset selection phase. We first present a unified formatting and labeling schema for all 36 claims datasets and 9 paragraph datasets. Next, we establish state-of-the-art baselines using GPT-4 with and without web search to collect evidence for veracity. Following this analysis, we find that standard evaluation metrics like accuracy and F1, when computed simply in relation to ground truth labels, are no longer sufficient to evaluate leading generative methods for misinformation detection and could lead to invalid conclusions. To address this challenge, we present and test an alternative evaluation approach, emphasizing the need for future research in this domain.
Finally, our findings come together in our recommendation for a new practice of Evaluation Quality Assessment (EQA), which offers a tool to improve data selection and evaluation methodologies. In summary, we present a guide to misinformation datasets, including:
  • The largest scale collection of misinformation datasets, CDL-MD, with a unified labeling schema, made easily accessible through a HuggingFace Repository; along with baseline performances on these datasets.
  • An essential toolkit, called CDL-DQA, for quality evaluation of misinformation datasets to analyze spurious correlations as well as the feasibility of the datapoints. Applying this toolkit to CDL-MD reveals numerous quality issues (summarized in Table 1), such as how the majority of claims datasets contain numerous examples whose veracity may be impossible to evaluate.
  • Recommended practices of EQA, for research proposing misinformation detection methods. Informed by our findings of severe quality issues in both data and evaluation procedures, such as how simple metrics like accuracy and F1 may be obsolete in this domain, we propose that assessment of evaluation quality is an indispensable component for future research here.